OpenStudy (anonymous):

f(x) + f(x+4) = f(x+2) + f(x+6) Find the period of the real valued function satisfying this equation

OpenStudy (anonymous):

\[ f(x) + f(x+4) = f(x+2) + f(x+6) \]Likewise, if we plug \(x+2\) into \(x\) we get::\[ f((x+2)) + f((x+2)+4) = f((x+2)+2) + f((x+2)+6) \]This simplifies to: \[ f(x+2) + f(x+6) = f(x+4) + f(x+8) \]This part is interesting because it show us: \[ f(x) + f(x+4) = f(x+2) + f(x+6) = f(x+4) + f(x+8) \]Meaning: \[ f(x) + f(x+4) = f(x+4) + f(x+8) \]Now if we subtract both sides by \(f(x+4)\), then we get: \[ f(x) = f(x+8) \]So at the very least, we have a period of \(8\).

OpenStudy (anonymous):

thank you